# VisJudge-Bench: Aesthetics and Quality Assessment of Visualizations

### v2

@@ C.1.1 Single Visualization Evaluation Criteria
Prompt: Single Visualization - Data Fidelity Evaluation Questions and Scoring Criteria (1-5 Scale) Description: Evaluates whether the visual encodings faithfully represent the data without misleading distortions. Focus on: appropriate axes and baselines (e.g., zero baseline for bar charts), reasonable scale ranges and tick intervals, consistency between encodings (position/length/angle/area/color) and numeric labels, and absence of aspect-ratio deformation, cropping/stretching, 3D distortion, or improper broken axes. Scoring criteria: S 1. Severe misrepresentation or impossible to judge due to missing/unreadable key elements (titles, legends, axes/ticks/labels, units, data labels, baselines, etc.), or clear distortions (e.g., non-zero baseline in bars exaggerating differences, 3D effects causing misread, obvious aspect-ratio deformation). S 2. Important issues likely to mislead: inappropriate baseline/scale range, inconsistent encoding vs labels, selective ranges that exaggerate differences, or partial key elements missing causing notable uncertainty. S 3. Basically acceptable with minor uncertainties: some elements may be missing or hard to read but no obvious distortion; scales and encodings are mostly reasonable and roughly consistent with labels. S 4. Accurate and standardized: axes/baselines are appropriate, scale ranges and ticks are reasonable, encodings match labels, and no visible distortion or manipulation is present. S 5. Highly faithful representation: all relevant elements are present and clear; axes/baselines and scale choices are well-justified (e.g., proper zero baseline for bars, appropriate linear/log choice), encodings and labels are fully consistent, with no cropping, deformation, or 3D misuse.
@@
Prompt: Single Visualization - Semantic Readability Evaluation Questions and Scoring Criteria (1-5 Scale) Description: Based on complete and clear chart elements, evaluates whether users can understand the meaning of these elements and the information the chart conveys, including clarity of visual encoding (whether meanings of colors, shapes, sizes are clear) and clarity of information communication (whether users can accurately understand chart content). Scoring criteria: S 1. Chart element meanings are completely unclear, cannot understand what visual encoding represents, users cannot obtain meaningful information. S 2. Chart element meanings are vague and unclear, insufficient visual encoding explanations, users have difficulty understanding and need extensive guessing to understand chart content. S 3. Chart element meanings are basically clear, users can understand main information, but still have understanding barriers in some encoding or labeling aspects. S 4. Chart element meanings are clear and definite, visual encoding has appropriate explanations, users can smoothly understand the information conveyed by the chart. S 5. Chart element meanings are completely clear, visual encoding explanations are detailed and accurate, users can understand all chart information without any barriers.
@@
Prompt: Single Visualization - Insight Discovery Evaluation Questions and Scoring Criteria (1-5 Scale) Description: Evaluates whether the chart can easily identify significant and meaningful data insights, trends, patterns, or anomalies that must have actual value and guiding significance for users, rather than trivial or obvious general information. Scoring criteria: S 1. Completely unable to provide any valuable insights, chaotic information presentation, no significant or meaningful discoveries, insight delivery completely failed. S 2. Difficult to identify meaningful insights, most presented information is trivial or obvious, lacking practical value, insight discovery difficult. S 3. Can identify some basic data patterns or trends, but the significance and practicality of insights are average, with limited guiding value for users. S 4. Can clearly identify obvious and useful data insights and trends, discovered patterns or anomalies have certain practical significance and reference value. S 5. Very intuitively reveals profound and highly significant data insights, can discover significant trend changes, important anomalies, or key patterns that have important guiding value for decision-making or understanding.
@@
Prompt: Single Visualization - Design Style Evaluation Questions and Scoring Criteria (1-5 Scale) Description: Evaluates whether the chart design has innovation and uniqueness, whether it can stand out from many visualization works through novel design elements and unique style characteristics. Scoring criteria: S 1. Design completely lacks innovation, style is outdated or shows obvious imitation, no uniqueness or originality. S 2. Design lacks innovation, style is quite ordinary, basically uses common design techniques, insufficient uniqueness. S 3. Design has some innovative elements, but overall is relatively conventional, uniqueness not prominent enough, innovation level is average. S 4. Design has strong innovation, style is quite unique, has novel design elements, with certain originality. S 5. Design is extremely innovative, style is unique and novel, uses innovative design elements or expression techniques, with strong originality and recognizability.
@@
Prompt: Single Visualization - Visual Composition Evaluation Questions and Scoring Criteria (1-5 Scale) Description: Evaluates whether the spatial layout of chart elements is reasonable, including whether element positions, size proportions, spacing distribution and other spatial relationships are appropriately and orderly arranged. Scoring criteria: S 1. Spatial layout is seriously unreasonable, element positions are chaotic, proportions severely unbalanced, spacing distribution is disorderly. S 2. Spatial layout has obvious unreasonable aspects, improper element positions, unbalanced proportions, uneven spacing distribution. S 3. Spatial layout is basically reasonable, element positions are acceptable, proportional relationships are basically coordinated, but may have small layout issues. S 4. Spatial layout is quite reasonable, element positions are appropriate, good proportional relationships, reasonable spacing distribution, proper space utilization. S 5. Spatial layout is very reasonable, element positions are perfect, proportions are coordinated, spacing distribution is even, space utilization efficiency is extremely high.
@@
Prompt: Single Visualization - Color Harmony Evaluation Questions and Scoring Criteria (1-5 Scale) Description: Evaluates whether chart color choices are coordinated, contrast is appropriate, color quantity is reasonable, and whether it avoids too many colors causing visual fatigue. Scoring criteria: S 1. Color choices are chaotic, extensive use of uncoordinated colors, serious visual fatigue, completely unable to effectively guide user understanding, color use severely affects information delivery. S 2. Color choices are not coordinated enough, using too many colors (5 or more), inappropriate contrast, causing obvious visual fatigue and understanding difficulties. S 3. Color choices are basically coordinated, but color quantity is excessive (4-5 colors), may cause slight visual complexity, somewhat affecting understanding. S 4. Color choices are relatively coordinated, appropriate contrast, reasonable color quantity control (3-4 colors), basically avoiding visual confusion, well guiding user understanding. S 5. Color choices are very coordinated, appropriate contrast, limited color scheme (usually 1-3 main colors), colors are simple and unified, completely avoiding visual fatigue, effectively maintaining user attention focus.
@@ C.1.2 Multiple Visualization Evaluation Criteria
Prompt: Multiple Visualization - Data Fidelity Evaluation Questions and Scoring Criteria (1-5 Scale) Description: Evaluates whether multiple coordinated views faithfully represent data without distortion individually and in combination. Check appropriate axes/baselines, reasonable scale ranges, consistency between encodings and labels, and cross-view consistency (e.g., comparable scales where appropriate), avoiding aspect-ratio deformation, cropping, 3D misuse, or improper broken axes. Scoring criteria: S 1. Severe misrepresentation or impossible to judge due to missing/unreadable key elements in one or more views, or clear distortions (non-zero bar baseline where required, 3D effects, obvious deformation), leading to unreliable reading. S 2. Important issues exist in one or more views (inappropriate baseline/scale, inconsistent encoding vs labels, selective ranges), or partial key elements missing causing notable cross-view uncertainty. S 3. Basically acceptable overall: minor uncertainties or small issues in some views, but no obvious distortion; scales and encodings are mostly reasonable across views. S 4. Accurate and standardized across views: appropriate axes/baselines and scales; encodings match labels; no visible distortion; reasonable cross-view scale alignment where needed. S 5. Highly faithful multi-vis representation: all views present complete, clear elements; well-justified axes/baselines and scale choices; full consistency between encodings and labels; no cropping/deformation/3D misuse; coherent cross-view comparability.
@@
Prompt: Multiple Visualization - Semantic Readability Evaluation Questions and Scoring Criteria (1-5 Scale) Description: Based on complete and clear chart elements, evaluates whether users can understand the meaning of these elements and the information the chart conveys, including clarity of visual encoding (whether meanings of colors, shapes, sizes are clear) and clarity of information communication (whether users can accurately understand chart content). Scoring criteria: S 1. Chart element meanings are completely unclear, cannot understand what visual encoding represents, users cannot obtain meaningful information. S 2. Chart element meanings are vague and unclear, insufficient visual encoding explanations, users have difficulty understanding and need extensive guessing to understand chart content. S 3. Chart element meanings are basically clear, users can understand main information, but still have understanding barriers in some encoding or labeling aspects. S 4. Chart element meanings are clear and definite, visual encoding has appropriate explanations, users can smoothly understand the information conveyed by the chart. S 5. Chart element meanings are completely clear, visual encoding explanations are detailed and accurate, users can understand all chart information without any barriers.
@@
Prompt: Multiple Visualization - Insight Discovery Evaluation Questions and Scoring Criteria (1-5 Scale) Description: Evaluates whether the entire multi-vis visualization can easily identify significant and meaningful data insights, and whether multiple views synergize to reveal deeper and more valuable comprehensive insights than single views. Scoring criteria: S 1. Completely unable to provide any valuable comprehensive insights, lack of coordination between views, no significant or meaningful discoveries, insight delivery completely failed. S 2. Difficult to identify meaningful comprehensive insights, most information presented by multiple views is trivial or independent, lacking synergistic value, insight discovery difficult. S 3. Can identify some basic multi-vis insights, but the significance and practicality of insights are average, limited cooperation between views, guiding value not prominent enough. S 4. Can clearly identify obvious and useful multi-dimensional insights, good cooperation between views, discovered comprehensive patterns have certain practical significance and reference value. S 5. Very intuitively reveals profound and highly significant comprehensive insights, perfect synergy between multiple views, can discover significant cross-dimensional patterns, important correlations, or key anomalies with important guiding value.
@@
Prompt: Multiple Visualization - Design Style Evaluation Questions and Scoring Criteria (1-5 Scale) Description: Evaluates whether the overall design of multi-vis visualization has innovation and uniqueness, whether it can stand out from many visualization works through novel design elements, unique style characteristics, and coordinated style language. Scoring criteria: S 1. Multi-vis design completely lacks innovation, style is outdated or views have chaotic styles, no uniqueness or originality. S 2. Multi-vis design lacks innovation, style is quite ordinary, basically uses common design techniques, insufficient uniqueness. S 3. Multi-vis design has some innovative elements, but overall is relatively conventional, style characteristics not prominent enough, innovation level is average. S 4. Multi-vis design has strong innovation, style is quite unique, coordinated style among views with certain originality. S 5. Multi-vis design is extremely innovative, style is unique and novel, views maintain unified style while showing strong originality and recognizability.
@@
Prompt: Multiple Visualization - Visual Composition Evaluation Questions and Scoring Criteria (1-5 Scale) Description: Evaluates whether the spatial layout of individual sub-views and the overall arrangement of multiple views are reasonable, including element composition within individual views, size proportions between multiple views, arrangement methods, spacing distribution, etc., forming a harmonious and orderly visual whole. Scoring criteria: S 1. Sub-view layouts are chaotic or overall multi-vis arrangement is seriously unreasonable, size proportions severely unbalanced, spacing distribution is disorderly. S 2. Sub-view layouts or multi-vis arrangement have obvious unreasonable aspects, inappropriate size proportions, uneven spacing distribution. S 3. Sub-view layouts are basically reasonable, overall multi-vis arrangement is acceptable, but may have small issues in proportional relationships or spacing handling. S 4. Sub-view layouts are good, overall multi-vis arrangement is reasonable, appropriate size proportions, proper spacing distribution, good space utilization. S 5. Sub-view layouts are perfect, overall multi-vis arrangement is extremely reasonable, coordinated size proportions, even spacing distribution, extremely high space utilization efficiency.
@@
Prompt: Multiple Visualization - Color Harmony Evaluation Questions and Scoring Criteria (1-5 Scale) Description: Evaluates whether color choices in multi-vis are appropriate, including color combinations within each sub-view, coordination of color schemes between multiple views, unity of overall tone, and appropriate control of color quantity and saturation. Scoring criteria: S 1. Sub-view color choices are very inappropriate, overall color matching among multiple views is chaotic, tone conflicts or color use is seriously inappropriate. S 2. Sub-view color choices are not appropriate enough, overall color matching among multiple views is not coordinated enough, tone not unified or color use inappropriate. S 3. Sub-view color choices are basically appropriate, overall color matching among multiple views is acceptable, but may have small issues in tone unity or saturation. S 4. Sub-view color choices are relatively appropriate, overall color schemes among multiple views are quite coordinated, tone basically unified, appropriate color use. S 5. Sub-view color choices are very appropriate, overall color schemes among multiple views are highly coordinated and unified, perfect tone matching, appropriate control of color quantity and saturation.

### v3

@@ C.1.1 Single Visualization Evaluation Criteria
Prompt: Single Visualization - Data Fidelity Evaluation Questions and Scoring Criteria (1–5 Scale) Description: Evaluates whether the visual encodings faithfully represent the data without misleading distortions. Focus on: appropriate axes and baselines (e.g., zero baseline for bar charts), reasonable scale ranges and tick intervals, consistency between encodings (position/length/angle/area/color) and numeric labels, and absence of aspect-ratio deformation, cropping/stretching, 3D distortion, or improper broken axes. Scoring criteria: S 1. Severe misrepresentation or impossible to judge due to missing/unreadable key elements (titles, legends, axes/ticks/labels, units, data labels, baselines, etc.), or clear distortions (e.g., non-zero baseline in bars exaggerating differences, 3D effects causing misread, obvious aspect-ratio deformation). S 2. Important issues likely to mislead: inappropriate baseline/scale range, inconsistent encoding vs labels, selective ranges that exaggerate differences, or partial key elements missing causing notable uncertainty. S 3. Basically acceptable with minor uncertainties: some elements may be missing or hard to read but no obvious distortion; scales and encodings are mostly reasonable and roughly consistent with labels. S 4. Accurate and standardized: axes/baselines are appropriate, scale ranges and ticks are reasonable, encodings match labels, and no visible distortion or manipulation is present. S 5. Highly faithful representation: all relevant elements are present and clear; axes/baselines and scale choices are well-justified (e.g., proper zero baseline for bars, appropriate linear/log choice), encodings and labels are fully consistent, with no cropping, deformation, or 3D misuse.
@@
Prompt: Single Visualization - Semantic Readability Evaluation Questions and Scoring Criteria (1–5 Scale) Description: Based on complete and clear chart elements, evaluates whether users can understand the meaning of these elements and the information the chart conveys, including clarity of visual encoding (whether meanings of colors, shapes, sizes are clear) and clarity of information communication (whether users can accurately understand chart content). Scoring criteria: S 1. Chart element meanings are completely unclear, cannot understand what visual encoding represents, users cannot obtain meaningful information. S 2. Chart element meanings are vague and unclear, insufficient visual encoding explanations, users have difficulty understanding and need extensive guessing to understand chart content. S 3. Chart element meanings are basically clear, users can understand main information, but still have understanding barriers in some encoding or labeling aspects. S 4. Chart element meanings are clear and definite, visual encoding has appropriate explanations, users can smoothly understand the information conveyed by the chart. S 5. Chart element meanings are completely clear, visual encoding explanations are detailed and accurate, users can understand all chart information without any barriers.
@@
Prompt: Single Visualization - Insight Discovery Evaluation Questions and Scoring Criteria (1–5 Scale) Description: Evaluates whether the chart can easily identify significant and meaningful data insights, trends, patterns, or anomalies that must have actual value and guiding significance for users, rather than trivial or obvious general information. Scoring criteria: S 1. Completely unable to provide any valuable insights, chaotic information presentation, no significant or meaningful discoveries, insight delivery completely failed. S 2. Difficult to identify meaningful insights, most presented information is trivial or obvious, lacking practical value, insight discovery difficult. S 3. Can identify some basic data patterns or trends, but the significance and practicality of insights are average, with limited guiding value for users. S 4. Can clearly identify obvious and useful data insights and trends, discovered patterns or anomalies have certain practical significance and reference value. S 5. Very intuitively reveals profound and highly significant data insights, can discover significant trend changes, important anomalies, or key patterns that have important guiding value for decision-making or understanding.
@@
Prompt: Single Visualization - Design Style Evaluation Questions and Scoring Criteria (1–5 Scale) Description: Evaluates whether the chart design has innovation and uniqueness, whether it can stand out from many visualization works through novel design elements and unique style characteristics. Scoring criteria: S 1. Design completely lacks innovation, style is outdated or shows obvious imitation, no uniqueness or originality. S 2. Design lacks innovation, style is quite ordinary, basically uses common design techniques, insufficient uniqueness. S 3. Design has some innovative elements, but overall is relatively conventional, uniqueness not prominent enough, innovation level is average. S 4. Design has strong innovation, style is quite unique, has novel design elements, with certain originality. S 5. Design is extremely innovative, style is unique and novel, uses innovative design elements or expression techniques, with strong originality and recognizability.
@@
Prompt: Single Visualization - Visual Composition Evaluation Questions and Scoring Criteria (1–5 Scale) Description: Evaluates whether the spatial layout of chart elements is reasonable, including whether element positions, size proportions, spacing distribution and other spatial relationships are appropriately and orderly arranged. Scoring criteria: S 1. Spatial layout is seriously unreasonable, element positions are chaotic, proportions severely unbalanced, spacing distribution is disorderly. S 2. Spatial layout has obvious unreasonable aspects, improper element positions, unbalanced proportions, uneven spacing distribution. S 3. Spatial layout is basically reasonable, element positions are acceptable, proportional relationships are basically coordinated, but may have small layout issues. S 4. Spatial layout is quite reasonable, element positions are appropriate, good proportional relationships, reasonable spacing distribution, proper space utilization. S 5. Spatial layout is very reasonable, element positions are perfect, proportions are coordinated, spacing distribution is even, space utilization efficiency is extremely high.
@@
Prompt: Single Visualization - Color Harmony Evaluation Questions and Scoring Criteria (1–5 Scale) Description: Evaluates whether chart color choices are coordinated, contrast is appropriate, color quantity is reasonable, and whether it avoids too many colors causing visual fatigue. Scoring criteria: S 1. Color choices are chaotic, extensive use of uncoordinated colors, serious visual fatigue, completely unable to effectively guide user understanding, color use severely affects information delivery. S 2. Color choices are not coordinated enough, using too many colors (5 or more), inappropriate contrast, causing obvious visual fatigue and understanding difficulties. S 3. Color choices are basically coordinated, but color quantity is excessive (4–5 colors), may cause slight visual complexity, somewhat affecting understanding. S 4. Color choices are relatively coordinated, appropriate contrast, reasonable color quantity control (3–4 colors), basically avoiding visual confusion, well guiding user understanding. S 5. Color choices are very coordinated, appropriate contrast, limited color scheme (usually 1–3 main colors), colors are simple and unified, completely avoiding visual fatigue, effectively maintaining user attention focus.
@@ C.1.2 Multiple Visualization Evaluation Criteria
Prompt: Multiple Visualization - Data Fidelity Evaluation Questions and Scoring Criteria (1–5 Scale) Description: Evaluates whether multiple coordinated views faithfully represent data without distortion individually and in combination. Check appropriate axes/baselines, reasonable scale ranges, consistency between encodings and labels, and cross-view consistency (e.g., comparable scales where appropriate), avoiding aspect-ratio deformation, cropping, 3D misuse, or improper broken axes. Scoring criteria: S 1. Severe misrepresentation or impossible to judge due to missing/unreadable key elements in one or more views, or clear distortions (non-zero bar baseline where required, 3D effects, obvious deformation), leading to unreliable reading. S 2. Important issues exist in one or more views (inappropriate baseline/scale, inconsistent encoding vs labels, selective ranges), or partial key elements missing causing notable cross-view uncertainty. S 3. Basically acceptable overall: minor uncertainties or small issues in some views, but no obvious distortion; scales and encodings are mostly reasonable across views. S 4. Accurate and standardized across views: appropriate axes/baselines and scales; encodings match labels; no visible distortion; reasonable cross-view scale alignment where needed. S 5. Highly faithful multi-vis representation: all views present complete, clear elements; well-justified axes/baselines and scale choices; full consistency between encodings and labels; no cropping/deformation/3D misuse; coherent cross-view comparability.
@@
Prompt: Multiple Visualization - Semantic Readability Evaluation Questions and Scoring Criteria (1–5 Scale) Description: Based on complete and clear chart elements, evaluates whether users can understand the meaning of these elements and the information the chart conveys, including clarity of visual encoding (whether meanings of colors, shapes, sizes are clear) and clarity of information communication (whether users can accurately understand chart content). Scoring criteria: S 1. Chart element meanings are completely unclear, cannot understand what visual encoding represents, users cannot obtain meaningful information. S 2. Chart element meanings are vague and unclear, insufficient visual encoding explanations, users have difficulty understanding and need extensive guessing to understand chart content. S 3. Chart element meanings are basically clear, users can understand main information, but still have understanding barriers in some encoding or labeling aspects. S 4. Chart element meanings are clear and definite, visual encoding has appropriate explanations, users can smoothly understand the information conveyed by the chart. S 5. Chart element meanings are completely clear, visual encoding explanations are detailed and accurate, users can understand all chart information without any barriers.
@@
Prompt: Multiple Visualization - Insight Discovery Evaluation Questions and Scoring Criteria (1–5 Scale) Description: Evaluates whether the entire multi-vis visualization can easily identify significant and meaningful data insights, and whether multiple views synergize to reveal deeper and more valuable comprehensive insights than single views. Scoring criteria: S 1. Completely unable to provide any valuable comprehensive insights, lack of coordination between views, no significant or meaningful discoveries, insight delivery completely failed. S 2. Difficult to identify meaningful comprehensive insights, most information presented by multiple views is trivial or independent, lacking synergistic value, insight discovery difficult. S 3. Can identify some basic multi-vis insights, but the significance and practicality of insights are average, limited cooperation between views, guiding value not prominent enough. S 4. Can clearly identify obvious and useful multi-dimensional insights, good cooperation between views, discovered comprehensive patterns have certain practical significance and reference value. S 5. Very intuitively reveals profound and highly significant comprehensive insights, perfect synergy between multiple views, can discover significant cross-dimensional patterns, important correlations, or key anomalies with important guiding value.
@@
Prompt: Multiple Visualization - Design Style Evaluation Questions and Scoring Criteria (1–5 Scale) Description: Evaluates whether the overall design of multi-vis visualization has innovation and uniqueness, whether it can stand out from many visualization works through novel design elements, unique style characteristics, and coordinated style language. Scoring criteria: S 1. Multi-vis design completely lacks innovation, style is outdated or views have chaotic styles, no uniqueness or originality. S 2. Multi-vis design lacks innovation, style is quite ordinary, basically uses common design techniques, insufficient uniqueness. S 3. Multi-vis design has some innovative elements, but overall is relatively conventional, style characteristics not prominent enough, innovation level is average. S 4. Multi-vis design has strong innovation, style is quite unique, coordinated style among views with certain originality. S 5. Multi-vis design is extremely innovative, style is unique and novel, views maintain unified style while showing strong originality and recognizability.
@@
Prompt: Multiple Visualization - Visual Composition Evaluation Questions and Scoring Criteria (1–5 Scale) Description: Evaluates whether the spatial layout of individual sub-views and the overall arrangement of multiple views are reasonable, including element composition within individual views, size proportions between multiple views, arrangement methods, spacing distribution, etc., forming a harmonious and orderly visual whole. Scoring criteria: S 1. Sub-view layouts are chaotic or overall multi-vis arrangement is seriously unreasonable, size proportions severely unbalanced, spacing distribution is disorderly. S 2. Sub-view layouts or multi-vis arrangement have obvious unreasonable aspects, inappropriate size proportions, uneven spacing distribution. S 3. Sub-view layouts are basically reasonable, overall multi-vis arrangement is acceptable, but may have small issues in proportional relationships or spacing handling. S 4. Sub-view layouts are good, overall multi-vis arrangement is reasonable, appropriate size proportions, proper spacing distribution, good space utilization. S 5. Sub-view layouts are perfect, overall multi-vis arrangement is extremely reasonable, coordinated size proportions, even spacing distribution, extremely high space utilization efficiency.
@@
Prompt: Multiple Visualization - Color Harmony Evaluation Questions and Scoring Criteria (1–5 Scale) Description: Evaluates whether color choices in multi-vis are appropriate, including color combinations within each sub-view, coordination of color schemes between multiple views, unity of overall tone, and appropriate control of color quantity and saturation. Scoring criteria: S 1. Sub-view color choices are very inappropriate, overall color matching among multiple views is chaotic, tone conflicts or color use is seriously inappropriate. S 2. Sub-view color choices are not appropriate enough, overall color matching among multiple views is not coordinated enough, tone not unified or color use inappropriate. S 3. Sub-view color choices are basically appropriate, overall color matching among multiple views is acceptable, but may have small issues in tone unity or saturation. S 4. Sub-view color choices are relatively appropriate, overall color schemes among multiple views are quite coordinated, tone basically unified, appropriate color use. S 5. Sub-view color choices are very appropriate, overall color schemes among multiple views are highly coordinated and unified, perfect tone matching, appropriate control of color quantity and saturation.